H. G. Wells

                                                                 9600 words
                                                            Disposable Copy

                          THE COUNTRY OF THE BLIND
                              by H. G. Wells
     Three hundred miles and more from Chimborazo, one hundred from the
snows of Cotopaxi, in the wildest wastes of Ecuador's Andes, there lies that
mysterious mountain valley, cut off from all the world of men, the Country
of the Blind.  Long years ago that valley lay so far open to the world that
men might come at last through frightful gorges and over an icy pass into
its equable meadows, and thither indeed men came, a family or so of Peruvian
half-breeds fleeing from the lust and tyranny of an evil Spanish ruler.
Then came the stupendous outbreak of Mindobamba, when it was night in Quito
for seventeen days, and the water was boiling at Yaguachi and all the fish
floating dying even as far as Guayaquil; everywhere along the Pacific slopes

                                                            Wells/BLIND/2
there were land-slips and swift thawings and sudden floods, and one whole
side of the old Arauca crest slipped and came down in thunder, and cut off
the Country of the Blind for ever from the exploring feet of men.  But one
of these early settlers had chanced to be on the hither side of the gorges
when the world had so terribly shaken itself, and he perforce had to forget
his wife and his child and all the friends and possessions he had left up
there, and start life over again in the lower world.  He started it again
but ill, blindness overtook him, and he died of punishment in the mines; but
the story he told begot a legend that lingers along the length of the
Cordilleras of the Andes to this day.
     He told of his reason for venturing back from that fastness, into which
he had first been carried lashed to a llama, beside a vast bale of gear,
when he was a child.  The valley, he said, had in it all that the heart of
man could desire–-sweet water, pasture, an even climate, slopes of rich
brown soil with tangles of a shrub that bore an excellent fruit, and on one
side great hanging forests of pine that held the avalanches high.  Far
overhead, on three sides, vast cliffs of grey-green rock were capped by
cliffs of ice; but the glacier stream came not to them, but flowed away by
the farther slopes, and only now and then huge ice masses fell on the valley
side.  In this valley it neither rained nor snowed, but the abundant springs
gave a rich green pasture, that irrigation would spread over all the valley
space.  The settlers did well indeed there.  Their beasts did well and
multiplied, and but one thing marred their happiness.  Yet it was enough to
mar it greatly.  A strange disease had come upon them and had made all the
children born to them there–-and, indeed, several older children

                                                            Wells/BLIND/3
also–-blind.  It was to seek some charm or antidote against this plague of
blindness that he had with fatigue and danger and difficulty returned down
the gorge.  In those days, in such cases, men did not think of germs and
infections, but of sins, and it seemed to him that the reason of this
affliction must he in the negligence of these priestless immigrants to set
up a shrine so soon as they entered the valley.  He wanted a shrine–-a
handsome, cheap, effectual shrine–-to be erected in the valley; he wanted
relics and such-like potent things of faith, blessed objects and mysterious
medals and prayers.  In his wallet he had a bar of native silver for which
he would not account; he insisted there was none in the valley with
something of the insistence of an inexpert liar.  They had all clubbed their
money and ornaments together, having little need for such treasure up there,
he said, to buy them holy help against their ill.  I figure this dim-eyed
young mountaineer, sunburnt, gaunt, and anxious, hat brim clutched
feverishly, a man all unused to the ways of the lower world, telling this
story to some keen-eyed, attentive priest before the great convulsion; I can
picture him presently seeking to return with pious and infallible remedies
against that trouble, and the infinite dismay with which he must have faced
the tumbled vastness where the gorge had once come out.  But the rest of his
story of mischances is lost to me, save that I know of his evil death after
several years.  Poor stray from that remoteness!  The stream that had once
made the gorge now bursts from the mouth of a rocky cave, and the legend his
poor, ill-told story set going developed into the legend of a race of blind
men somewhere "over there" one may still hear to-day.
     And amidst the little population of that now isolated and forgotten

                                                            Wells/BLIND/4
valley the disease ran its course.  The old became groping, the young saw
but dimly, and the children that were born to them never saw at all.  But
life was very easy in that snow-rimmed basin, lost to all the world, with
neither thorns nor briers, with no evil insects nor any beasts save the
gentle breed of llamas they had lugged and thrust and followed up the beds
of the shrunken rivers in the gorges up which they had come.  The seeing had
become purblind so gradually that they scarcely noticed their loss.  They
guided the sightless youngsters hither and thither until they knew the whole
valley marvellously, and when at last sight died out among them the race
lived on.  They had even time to adapt themselves to the blind control of
fire, which they made carefully in stoves of stone.  They were a simple
strain of people at the first, unlettered, only slightly touched with the
Spanish civilisation, but with something of a tradition of the arts of old
Peru and of its lost philosophy.  Generation followed generation.  They
forgot many things; they devised many things.  Their tradition of the
greater world they came from became mythical in colour and uncertain.  In
all things save sight they were strong and able, and presently chance sent
one who had an original mind and who could talk and persuade among them, and
then afterwards another.  These two passed, leaving their effects, and the
little community grew in numbers and in understanding, and met and settled
social and economic problems that arose.  Generation followed generation.
Generation followed generation.  There came a time when a child was born who
was fifteen generations from that ancestor who went out of the valley with a
bar of silver to seek God's aid, and who never returned.  Thereabout it
chanced that a man came into this community from the outer world.  And this

                                                            Wells/BLIND/5
is the story of that man.
     He was a mountaineer from the country near Quito, a man who had been
down to the sea and had seen the world, a reader of books in an original
way, an acute and enterprising man, and he was taken on by a party of
Englishmen who had come out to Ecuador to climb mountains, to replace one of
their three Swiss guides who had fallen ill.  He climbed here and he climbed
there, and then came the attempt on Parascotopetl, the Matterhorn of the
Andes, in which he was lost to the outer world.  The story of that accident
has been written a dozen times.  Pointer's narrative is the best.  He tells
how the little party worked their difficult and almost vertical way up to
the very foot of the last and greatest precipice, and how they built a night
shelter amidst the snow upon a little shelf of rock, and, with a touch of
real dramatic power, how presently they found Nunez had gone from them.
They shouted, and there was no reply; shouted and whistled, and for the rest
of that night they slept no more.
     As the morning broke they saw the traces of his fall.  It seems
impossible he could have uttered a sound.  He had slipped eastward towards
the unknown side of the mountain; far below he had struck a steep slope of
snow, and ploughed his way down it in the midst of a snow avalanche.  His
track went straight to the edge of a frightful precipice, and beyond that
everything was hidden.  Far, far below, and hazy with distance, they could
see trees rising out of a narrow, shut-in valley–-the lost Country of the
Blind.  But they did not know it was the lost Country of the Blind, nor
distinguish it in any way from any other narrow streak of upland valley.
Unnerved by this disaster, they abandoned their attempt in the afternoon,

                                                            Wells/BLIND/6
and Pointer was called away to the war before he could make another attack.
To this day Parascotopetl lifts an unconquered crest, and Pointer's shelter
crumbles unvisited amidst the snows.
     And the man who fell survived.
     At the end of the slope he fell a thousand feet, and came down in the
midst of a cloud of snow upon a snow-slope even steeper than the one above.
Down this he was whirled, stunned and insensible, but without a bone broken
in his body; and then at last came to gentler slopes, and at last rolled out
and lay still, buried amidst a softening heap of the white masses that had
accompanied and saved him.  He came to himself with a dim fancy that he was
ill in bed; then realized his position with a mountaineer's intelligence and
worked himself loose and, after a rest or so, out until he saw the stars.
He rested flat upon his chest for a space, wondering where he was and what
had happened to him.  He explored his limbs, and discovered that several of
his buttons were gone and his coat turned over his head.  His knife had gone
from his pocket and his hat was lost, though he had tied it under his chin.
He recalled that he had been looking for loose stones to raise his piece of
the shelter wall.  His ice-axe had disappeared.
     He decided he must have fallen, and looked up to see, exaggerated by
the ghastly light of the rising moon, the tremendous flight he had taken.
For a while he lay, gazing blankly at the vast, pale cliff towering above,
rising moment by moment out of a subsiding tide of darkness.  Its
phantasmal, mysterious beauty held him for a space, and then he was seized
with a paroxysm of sobbing laughter .  .  .  .
     After a great interval of time he became aware that he was near the

                                                            Wells/BLIND/7
lower edge of the snow.  Below, down what was now a moon-lit and practicable
slope, he saw the dark and broken appearance of rock-strewn turf He
struggled to his feet, aching in every joint and limb, got down painfully
from the heaped loose snow about him, went downward until he was on the
turf, and there dropped rather than lay beside a boulder, drank deep from
the flask in his inner pocket, and instantly fell asleep .  .  .  .
     He was awakened by the singing of birds in the trees far below.
     He sat up and perceived he was on a little alp at the foot of a vast
precipice that sloped only a little in the gully down which he and his snow
had come.  Over against him another wall of rock reared itself against the
sky.  The gorge between these precipices ran east and west and was full of
the morning sunlight, which lit to the westward the mass of fallen mountain
that closed the descending gorge.  Below him it seemed there was a precipice
equally steep, but behind the snow in the gully he found a sort of
chimney-cleft dripping with snow-water, down which a desperate man might
venture.  He found it easier than it seemed, and came at last to another
desolate alp, and then after a rock climb of no particular difficulty, to a
steep slope of trees.  He took his bearings and turned his face up the
gorge, for he saw it opened out above upon green meadows, among which he now
glimpsed quite distinctly a cluster of stone huts of unfamiliar fashion.  At
times his progress was like clambering along the face of a wall, and after a
time the rising sun ceased to strike along the gorge, the voices of the
singing birds died away, and the air grew cold and dark about him.  But the
distant valley with its houses was all the brighter for that.  He came
presently to talus, and among the rocks he noted–-for he was an observant

                                                            Wells/BLIND/8
man–-an unfamiliar fern that seemed to clutch out of the crevices with
intense green hands.  He picked a frond or so and gnawed its stalk, and
found it helpful.
     About midday he came at last out of the throat of the gorge into the
plain and the sunlight.  He was stiff and weary; he sat down in the shadow
of a rock, filled up his flask with water from a spring and drank it down,
and remained for a time, resting before he went on to the houses.
     They were very strange to his eyes, and indeed the whole aspect of that
valley became, as he regarded it, queerer and more unfamiliar.  The greater
part of its surface was lush green meadow, starred with many beautiful
flowers, irrigated with extraordinary care, and bearing evidence of
systematic cropping piece by piece.  High up and ringing the valley about
was a wall, and what appeared to be a circumferential water channel, from
which the little trickles of water that fed the meadow plants came, and on
the higher slopes above this flocks of llamas cropped the scanty herbage.
Sheds, apparently shelters or feeding-places for the llamas, stood against
the boundary wall here and there.  The irrigation streams ran together into
a main channel down the centre of the valley, and this was enclosed on
either side by a wall breast high.  This gave a singularly urban quality to
this secluded place, a quality that was greatly enhanced by the fact that a
number of paths paved with black and white stones, and each with a curious
little kerb at the side, ran hither and thither in an orderly manner.  The
houses of the central village were quite unlike the casual and
higgledy-piggledy agglomeration of the mountain villages he knew; they stood
in a continuous row on either side of a central street of astonishing

                                                            Wells/BLIND/9
cleanness, here and there their parti-coloured facade was pierced by a door,
and not a solitary window broke their even frontage.  They were
parti-coloured with extraordinary irregularity, smeared with a sort of
plaster that was sometimes grey, sometimes drab, sometimes slate-coloured or
dark brown; and it was the sight of this wild plastering first brought the
word "blind" into the thoughts of the explorer.  "The good man who did
that," he thought, "must have been as blind as a bat."
     He descended a steep place, and so came to the wall and channel that
ran about the valley, near where the latter spouted out its surplus contents
into the deeps of the gorge in a thin and wavering thread of cascade.  He
could now see a number of men and women resting on piled heaps of grass, as
if taking a siesta, in the remoter part of the meadow, and nearer the
village a number of recumbent children, and then nearer at hand three men
carrying pails on yokes along a little path that ran from the encircling
wall towards the houses.  These latter were clad in garments of llama cloth
and boots and belts of leather, and they wore caps of cloth with back and
ear flaps.  They followed one another in single file, walking slowly and
yawning as they walked, like men who have been up all night.  There was
something so reassuringly prosperous and respectable in their bearing that
after a moment's hesitation Nunez stood forward as conspicuously as possible
upon his rock, and gave vent to a mighty shout that echoed round the valley.
     The three men stopped, and moved their heads as though they were
looking about them.  They turned their faces this way and that, and Nunez
gesticulated with freedom.  But they did not appear to see him for all his
gestures, and after a time, directing themselves towards the mountains far

                                                           Wells/BLIND/10
away to the right, they shouted as if in answer.  Nunez bawled again, and
then once more, and as he gestured ineffectually the word "blind" came up to
the top of his thoughts.  "The fools must be blind," he said.
     When at last, after much shouting and wrath, Nunez crossed the stream
by a little bridge, came through a gate in the wall, and approached them, he
was sure that they were blind.  He was sure that this was the Country of the
Blind of which the legends told.  Conviction had sprung upon him, and a
sense of great and rather enviable adventure.  The three stood side by side,
not looking at him, but with their ears directed towards him, judging him by
his unfamiliar steps.  They stood close together like men a little afraid,
and he could see their eyelids closed and sunken, as though the very balls
beneath had shrunk away.  There was an expression near awe on their faces.
     "A man," one said, in hardly recognisable Spanish.  "A man it is–-a man
or a spirit–-coming down from the rocks."
     But Nunez advanced with the confident steps of a youth who enters upon
life.  All the old stories of the lost valley and the Country of the Blind
had come back to his mind, and through his thoughts ran this old proverb, as
if it were a refrain:–-
     "In the Country of the Blind the One-Eyed Man is King."
     "In the Country of the Blind the One-Eyed Man is King."
     And very civilly he gave them greeting.  He talked to them and used his
eyes.
     "Where does he come from, brother Pedro?"  asked one.
     "Down out of the rocks."
     "Over the mountains I come," said Nunez, "out of the country beyond

                                                           Wells/BLIND/11
there–-where men can see.  From near Bogota–-where there are a hundred
thousands of people, and where the city passes out of sight."
     "Sight?"  muttered Pedro.  "Sight?"
     "He comes," said the second blind man, "out of the rocks."
     The cloth of their coats, Nunez saw was curious fashioned, each with a
different sort of stitching.
     They startled him by a simultaneous movement towards him, each with a
hand outstretched.  He stepped back from the advance of these spread
fingers.
     "Come hither," said the third blind man, following his motion and
clutching him neatly.
     And they held Nunez and felt him over, saying no word further until
they had done so.
     "Carefully," he cried, with a finger in his eye, and found they thought
that organ, with its fluttering lids, a queer thing in him.  They went over
it again.
     "A strange creature, Correa," said the one called Pedro.  "Feel the
coarseness of his hair.  Like a llama's hair."
     "Rough he is as the rocks that begot him," said Correa, investigating
Nunez's unshaven chin with a soft and slightly moist hand.  "Perhaps he will
grow finer."
     Nunez struggled a little under their examination, but they gripped him
firm.
     "Carefully," he said again.
     "He speaks," said the third man.  "Certainly he is a man."

                                                      Wells/BLIND/12
"Ugh!"  said Pedro, at the roughness of his coat.
"And you have come into the world?"  asked Pedro.
"OUT of the world.  Over mountains and glaciers; right over above
there, half-way to the sun.  Out of the great, big world that goes down,
twelve days' journey to the sea."
They scarcely seemed to heed him.  "Our fathers have told us men may be
made by the forces of Nature," said Correa.  "It is the warmth of things,
and moisture, and rottenness–-rottenness."
"Let us lead him to the elders," said Pedro.
"Shout first," said Correa, "lest the children be afraid.  This is a
marvellous occasion."
So they shouted, and Pedro went first and took Nunez by the hand to
lead him to the houses.
He drew his hand away.  "I can see," he said.
"See?"  said Correa.
"Yes; see," said Nunez, turning towards him, and stumbled against
Pedro's pail.
"His senses are still imperfect," said the third blind man.  "He
stumbles, and talks unmeaning words.  Lead him by the hand."
"As you will," said Nunez, and was led along laughing.
It seemed they knew nothing of sight.
Well, all in good time he would teach them.
He heard people shouting, and saw a number of figures gathering
together in the middle roadway of the village.
He found it tax his nerve and patience more than he had anticipated,

                                                           Wells/BLIND/13
that first encounter with the population of the Country of the Blind.  The
place seemed larger as he drew near to it, and the smeared plasterings
queerer, and a crowd of children and men and women (the women and girls he
was pleased to note had, some of them, quite sweet faces, for all that their
eyes were shut and sunken) came about him, holding on to him, touching him
with soft, sensitive hands, smelling at him, and listening at every word he
spoke.  Some of the maidens and children, however, kept aloof as if afraid,
and indeed his voice seemed coarse and rude beside their softer notes.  They
mobbed him.  His three guides kept close to him with an effect of
proprietorship, and said again and again, "A wild man out of the rocks."
     "Bogota," he said.  "Bogota.  Over the mountain crests."
     "A wild man–-using wild words," said Pedro.  "Did you hear
that–-"BOGOTA? His mind has hardly formed yet.  He has only the beginnings
of speech."
     A little boy nipped his hand.  "Bogota!"  he said mockingly.
     "Aye!  A city to your village.  I come from the great world–-where men
have eyes and see."
     "His name's Bogota," they said.
     "He stumbled," said Correa–-" stumbled twice as we came hither."
     "Bring him in to the elders."
     And they thrust him suddenly through a doorway into a room as black as
pitch, save at the end there faintly glowed a fire.  The crowd closed in
behind him and shut out all but the faintest glimmer of day, and before he
could arrest himself he had fallen headlong over the feet of a seated man.
His arm, outflung, struck the face of someone else as he went down; he felt

                                                           Wells/BLIND/14
the soft impact of features and heard a cry of anger, and for a moment he
struggled against a number of hands that clutched him.  It was a one-sided
fight.  An inkling of the situation came to him and he lay quiet.
     "I fell down," be said; I couldn't see in this pitchy darkness."
     There was a pause as if the unseen persons about him tried to
understand his words.  Then the voice of Correa said:  "He is but newly
formed.  He stumbles as he walks and mingles words that mean nothing with
his speech."
     Others also said things about him that he heard or understood
imperfectly.
     "May I sit up?"  he asked, in a pause.  "I will not struggle against
you again."
     They consulted and let him rise.
     The voice of an older man began to question him, and Nunez found
himself trying to explain the great world out of which he had fallen, and
the sky and mountains and such-like marvels, to these elders who sat in
darkness in the Country of the Blind.  And they would believe and understand
nothing whatever that he told them, a thing quite outside his expectation.
They would not even understand many of his words.  For fourteen generations
these people had been blind and cut off from all the seeing world; the names
for all the things of sight had faded and changed; the story of the outer
world was faded and changed to a child's story; and they had ceased to
concern themselves with anything beyond the rocky slopes above their
circling wall.  Blind men of genius had arisen among them and questioned the
shreds of belief and tradition they had brought with them from their seeing

                                                           Wells/BLIND/15
days, and had dismissed all these things as idle fancies and replaced them
with new and saner explanations.  Much of their imagination had shrivelled
with their eyes, and they had made for themselves new imaginations with
their ever more sensitive ears and finger-tips.  Slowly Nunez realised this:
that his expectation of wonder and reverence at his origin and his gifts was
not to be borne out; and after his poor attempt to explain sight to them had
been set aside as the confused version of a new-made being describing the
marvels of his incoherent sensations, he subsided, a little dashed, into
listening to their instruction.  And the eldest of the blind men explained
to him life and philosophy and religion, how that the world (meaning their
valley) had been first an empty hollow in the rocks, and then had come first
inanimate things without the gift of touch, and llamas and a few other
creatures that had little sense, and then men, and at last angels, whom one
could hear singing and making fluttering sounds, but whom no one could touch
at all, which puzzled Nunez greatly until he thought of the birds.
     He went on to tell Nunez how this time had been divided into the warm
and the cold, which are the blind equivalents of day and night, and how it
was good to sleep in the warm and work during the cold, so that now, but for
his advent, the whole town of the blind would have been asleep.  He said
Nunez must have been specially created to learn and serve the wisdom they
had acquired, and that for all his mental incoherency and stumbling
behaviour he must have courage and do his best to learn, and at that all the
people in the door-way murmured encouragingly.  He said the night–-for the
blind call their day night–-was now far gone, and it behooved everyone to go
back to sleep.  He asked Nunez if he knew how to sleep, and Nunez said he

                                                           Wells/BLIND/16
did, but that before sleep he wanted food.  They brought him food, llama's
milk in a bowl and rough salted bread, and led him into a lonely place to
eat out of their hearing, and afterwards to slumber until the chill of the
mountain evening roused them to begin their day again.  But Nunez slumbered
not at all.
     Instead, he sat up in the place where they had left him, resting his
limbs and turning the unanticipated circumstances of his arrival over and
over in his mind.
     Every now and then he laughed, sometimes with amusement and sometimes
with indignation.
     "Unformed mind!"  he said.  "Got no senses yet!  They little know
they've been insulting their Heaven-sent King and master .  .  .  .  .
     "I see I must bring them to reason.
     "Let me think.
     "Let me think."
     He was still thinking when the sun set.
     Nunez had an eye for all beautiful things, and it seemed to him that
the glow upon the snow-fields and glaciers that rose about the valley on
every side was the most beautiful thing he had ever seen.  His eyes went
from that inaccessible glory to the village and irrigated fields, fast
sinking into the twilight, and suddenly a wave of emotion took him, and he
thanked God from the bottom of his heart that the power of sight had been
given him.
     He heard a voice calling to him from out of the village.
     "Yaho there, Bogota!  Come hither!"

                                                      Wells/BLIND/17
At that he stood up, smiling.  He would show these people once and for
all what sight would do for a man.  They would seek him, but not find him.
"You move not, Bogota," said the voice.
He laughed noiselessly and made two stealthy steps aside from the path.
"Trample not on the grass, Bogota; that is not allowed."
Nunez had scarcely heard the sound he made himself.  He stopped,
amazed.
The owner of the voice came running up the piebald path towards him.
He stepped back into the pathway.  "Here I am," he said.
"Why did you not come when I called you?"  said the blind man.  "Must
you be led like a child?  Cannot you hear the path as you walk?"
Nunez laughed.  "I can see it," he said.
"There is no such word as SEE," said the blind man, after a pause.
"Cease this folly and follow the sound of my feet."
Nunez followed, a little annoyed.
"My time will come," he said.
"You'll learn," the blind man answered.  "There is much to learn in the
world."
"Has no one told you, `In the Country of the Blind the One-Eyed Man is
King?'"
"What is blind?" asked the blind man, carelessly, over his shoulder.
Four days passed and the fifth found the King of the Blind still
incognito, as a clumsy and useless stranger among his subjects.
It was, he found, much more difficult to proclaim himself than he had
supposed, and in the meantime, while he meditated his coup d'etat, he did

                                                           Wells/BLIND/18
what he was told and learnt the manners and customs of the Country of the
Blind.  He found working and going about at night a particularly irksome
thing, and he decided that that should be the first thing he would change.
     They led a simple, laborious life, these people, with all the elements
of virtue and happiness as these things can be understood by men.  They
toiled, but not oppressively; they had food and clothing sufficient for
their needs; they had days and seasons of rest; they made much of music and
singing, and there was love among them and little children.  It was
marvellous with what confidence and precision they went about their ordered
world.  Everything, you see, had been made to fit their needs; each of the
radiating paths of the valley area had a constant angle to the others, and
was distinguished by a special notch upon its kerbing; all obstacles and
irregularities of path or meadow had long since been cleared away; all their
methods and procedure arose naturally from their special needs.  Their
senses had become marvellously acute; they could hear and judge the
slightest gesture of a man a dozen paces away–-could hear the very beating
of his heart.  Intonation had long replaced expression with them, and
touches gesture, and their work with hoe and spade and fork was as free and
confident as garden work can be.  Their sense of smell was extraordinarily
fine; they could distinguish individual differences as readily as a dog can,
and they went about the tending of llamas, who lived among the rocks above
and came to the wall for food and shelter, with ease and confidence.  It was
only when at last Nunez sought to assert himself that he found how easy and
confident their movements could be.
     He rebelled only after he had tried persuasion.

                                                           Wells/BLIND/19
     He tried at first on several occasions to tell them of sight.  "Look
you here, you people," he said.  "There are things you do not understand in
me."
     Once or twice one or two of them attended to him; they sat with faces
downcast and ears turned intelligently towards him, and he did his best to
tell them what it was to see.  Among his hearers was a girl, with eyelids
less red and sunken than the others, so that one could almost fancy she was
hiding eyes, whom especially he hoped to persuade.  He spoke of the beauties
of sight, of watching the mountains, of the sky and the sunrise, and they
heard him with amused incredulity that presently became condemnatory.  They
told him there were indeed no mountains at all, but that the end of the
rocks where the llamas grazed was indeed the end of the world; thence sprang
a cavernous roof of the universe, from which the dew and the avalanches
fell; and when he maintained stoutly the world had neither end nor roof such
as they supposed, they said his thoughts were wicked.  So far as he could
describe sky and clouds and stars to them it seemed to them a hideous void,
a terrible blankness in the place of the smooth roof to things in which they
believed–-it was an article of faith with them that the cavern roof was
exquisitely smooth to the touch.  He saw that in some manner he shocked
them, and gave up that aspect of the matter altogether, and tried to show
them the practical value of sight.  One morning he saw Pedro in the path
called Seventeen and coming towards the central houses, but still too far
off for hearing or scent, and he told them as much.  "In a little while," he
prophesied, "Pedro will be here."  An old man remarked that Pedro had no
business on path Seventeen, and then, as if in confirmation, that individual

                                                           Wells/BLIND/20
as he drew near turned and went transversely into path Ten, and so back with
nimble paces towards the outer wall.  They mocked Nunez when Pedro did not
arrive, and afterwards, when he asked Pedro questions to clear his
character, Pedro denied and outfaced him, and was afterwards hostile to him.
     Then he induced them to let him go a long way up the sloping meadows
towards the wall with one complaisant individual, and to him he promised to
describe all that happened among the houses.  He noted certain goings and
comings, but the things that really seemed to signify to these people
happened inside of or behind the windowless houses–-the only things they
took note of to test him by–-and of those he could see or tell nothing; and
it was after the failure of this attempt, and the ridicule they could not
repress, that he resorted to force.  He thought of seizing a spade and
suddenly smiting one or two of them to earth, and so in fair combat showing
the advantage of eyes.  He went so far with that resolution as to seize his
spade, and then he discovered a new thing about himself, and that was that
it was impossible for him to hit a blind man in cold blood.
     He hesitated, and found them all aware that he had snatched up the
spade.  They stood all alert, with their heads on one side, and bent ears
towards him for what he would do next.
     "Put that spade down," said one, and he felt a sort of helpless horror.
He came near obedience.
     Then he had thrust one backwards against a house wall, and fled past
him and out of the village.
     He went athwart one of their meadows, leaving a track of trampled grass
behind his feet, and presently sat down by the side of one of their ways.

                                                           Wells/BLIND/21
He felt something of the buoyancy that comes to all men in the beginning of
a fight, but more perplexity.  He began to realise that you cannot even
fight happily with creatures who stand upon a different mental basis to
yourself.  Far away he saw a number of men carrying spades and sticks come
out of the street of houses and advance in a spreading line along the
several paths towards him.  They advanced slowly, speaking frequently to one
another, and ever and again the whole cordon would halt and sniff the air
and listen.
     The first time they did this Nunez laughed.  But afterwards he did not
laugh.
     One struck his trail in the meadow grass and came stooping and feeling
his way along it.
     For five minutes he watched the slow extension of the cordon, and then
his vague disposition to do something forthwith became frantic.  He stood
up, went a pace or so towards the circumferential wall, turned, and went
back a little way.  There they all stood in a crescent, still and listening.
     He also stood still, gripping his spade very tightly in both hands.
Should he charge them?
     The pulse in his ears ran into the rhythm of "In the Country of the
Blind the One-Eyed Man is King."
     Should he charge them?
     He looked back at the high and unclimbable wall behind–-unclimbable
because of its smooth plastering, but withal pierced with many little doors
and at the approaching line of seekers.  Behind these others were now coming
out of the street of houses.

                                                           Wells/BLIND/22
     Should he charge them?
     "Bogota!"  called one.  "Bogota!  where are you?"
     He gripped his spade still tighter and advanced down the meadows
towards the place of habitations, and directly he moved they converged upon
him.  "I'll hit them if they touch me," he swore; "by Heaven, I will.  I'll
hit."  He called aloud, "Look here, I'm going to do what I like in this
valley!  Do you hear?  I'm going to do what I like and go where I like."
     They were moving in upon him quickly, groping, yet moving rapidly.  It
was like playing blind man's buff with everyone blindfolded except one.
"Get hold of him!"  cried one.  He found himself in the arc of a loose curve
of pursuers.  He felt suddenly he must be active and resolute.
     "You don't understand," he cried, in a voice that was meant to be great
and resolute, and which broke.  "You are blind and I can see.  Leave me
alone!"
     "Bogota!  Put down that spade and come off the grass!"
     The last order, grotesque in its urban familiarity, produced a gust of
anger.  "I'll hurt you," he said, sobbing with emotion.  "By Heaven, I'll
hurt you!  Leave me alone!"
     He began to run–-not knowing clearly where to run.  He ran from the
nearest blind man, because it was a horror to hit him.  He stopped, and then
made a dash to escape from their closing ranks.  He made for where a gap was
wide, and the men on either side, with a quick perception of the approach of
his paces, rushed in on one another.  He sprang forward, and then saw he
must be caught, and SWISH! the spade had struck.  He felt the soft thud of
hand and arm, and the man was down with a yell of pain, and he was through.

                                                           Wells/BLIND/23
     Through!  And then he was close to the street of houses again, and
blind men, whirling spades and stakes, were running with a reasoned
swiftness hither and thither.
     He heard steps behind him just in time, and found a tall man rushing
forward and swiping at the sound of him.  He lost his nerve, hurled his
spade a yard wide of this antagonist, and whirled about and fled, fairly
yelling as he dodged another.
     He was panic-stricken.  He ran furiously to and fro, dodging when there
was no need to dodge, and, in his anxiety to see on every side of him at
once, stumbling.  For a moment he was down and they heard his fall.  Far
away in the circumferential wall a little doorway looked like Heaven, and he
set off in a wild rush for it.  He did not even look round at his pursuers
until it was gained, and he had stumbled across the bridge, clambered a
little way among the rocks, to the surprise and dismay of a young llama, who
went leaping out of sight, and lay down sobbing for breath.
     And so his coup d'etat came to an end.
     He stayed outside the wall of the valley of the blind for two nights
and days without food or shelter, and meditated upon the Unexpected.  During
these meditations he repeated very frequently and always with a profounder
note of derision the exploded proverb:  "In the Country of the Blind the
One-Eyed Man is King."  He thought chiefly of ways of fighting and
conquering these people, and it grew clear that for him no practicable way
was possible.  He had no weapons, and now it would be hard to get one.
     The canker of civilisation had got to him even in Bogota, and he could
not find it in himself to go down and assassinate a blind man.  Of course,

                                                           Wells/BLIND/24
if he did that, he might then dictate terms on the threat of assassinating
them all.  But–-Sooner or later he must sleep!  .  .  .  .
     He tried also to find food among the pine trees, to be comfortable
under pine boughs while the frost fell at night, and–- with less
confidence–-to catch a llama by artifice in order to try to kill it–-perhaps
by hammering it with a stone–-and so finally, perhaps, to eat some of it.
But the llamas had a doubt of him and regarded him with distrustful brown
eyes and spat when he drew near.  Fear came on him the second day and fits
of shivering.  Finally he crawled down to the wall of the Country of the
Blind and tried to make his terms.  He crawled along by the stream,
shouting, until two blind men came out to the gate and talked to him.
     "I was mad," he said.  "But I was only newly made."
     They said that was better.
     He told them he was wiser now, and repented of all he had done.
     Then he wept without intention, for he was very weak and ill now, and
they took that as a favourable sign.
     They asked him if he still thought he could SEE."
     "No," he said.  "That was folly.  The word means nothing.  Less than
nothing!"
     They asked him what was overhead.
     "About ten times ten the height of a man there is a roof above the
world–-of rock–-and very, very smooth.  So smooth–-so beautifully smooth .
.  "He burst again into hysterical tears.  "Before you ask me any more, give
me some food or I shall die!"
     He expected dire punishments, but these blind people were capable of

                                                           Wells/BLIND/25
toleration.  They regarded his rebellion as but one more proof of his
general idiocy and inferiority, and after they had whipped him they
appointed him to do the simplest and heaviest work they had for anyone to
do, and he, seeing no other way of living, did submissively what he was
told.
     He was ill for some days and they nursed him kindly.  That refined his
submission.  But they insisted on his lying in the dark, and that was a
great misery.  And blind philosophers came and talked to him of the wicked
levity of his mind, and reproved him so impressively for his doubts about
the lid of rock that covered their cosmic casserole that he almost doubted
whether indeed he was not the victim of hallucination in not seeing it
overhead.
     So Nunez became a citizen of the Country of the Blind, and these people
ceased to be a generalised people and became individualities to him, and
familiar to him, while the world beyond the mountains became more and more
remote and unreal.  There was Yacob, his master, a kindly man when not
annoyed; there was Pedro, Yacob's nephew; and there was Medina-sarote, who
was the youngest daughter of Yacob.  She was little esteemed in the world of
the blind, because she had a clear-cut face and lacked that satisfying,
glossy smoothness that is the blind man's ideal of feminine beauty, but
Nunez thought her beautiful at first, and presently the most beautiful thing
in the whole creation.  Her closed eyelids were not sunken and red after the
common way of the valley, but lay as though they might open again at any
moment; and she had long eyelashes, which were considered a grave
disfigurement.  And her voice was weak and did not satisfy the acute hearing

                                                           Wells/BLIND/26
of the valley swains.  So that she had no lover.
     There came a time when Nunez thought that, could he win her, he would
be resigned to live in the valley for all the rest of his days.
     He watched her; he sought opportunities of doing her little services
and presently he found that she observed him.  Once at a rest-day gathering
they sat side by side in the dim starlight, and the music was sweet.  His
hand came upon hers and he dared to clasp it.  Then very tenderly she
returned his pressure.  And one day, as they were at their meal in the
darkness, he felt her hand very softly seeking him, and as it chanced the
fire leapt then, and he saw the tenderness of her face.
     He sought to speak to her.
     He went to her one day when she was sitting in the summer moonlight
spinning.  The light made her a thing of silver and mystery.  He sat down at
her feet and told her he loved her, and told her how beautiful she seemed to
him.  He had a lover's voice, he spoke with a tender reverence that came
near to awe, and she had never before been touched by adoration.  She made
him no definite answer, but it was clear his words pleased her.
     After that he talked to her whenever he could take an opportunity.  The
valley became the world for him, and the world beyond the mountains where
men lived by day seemed no more than a fairy tale he would some day pour
into her ears.  Very tentatively and timidly he spoke to her of sight.
     Sight seemed to her the most poetical of fancies, and she listened to
his description of the stars and the mountains and her own sweet white-lit
beauty as though it was a guilty indulgence.  She did not believe, she could
only half understand, but she was mysteriously delighted, and it seemed to

                                                           Wells/BLIND/27
him that she completely understood.
     His love lost its awe and took courage.  Presently he was for demanding
her of Yacob and the elders in marriage, but she became fearful and delayed.
And it was one of her elder sisters who first told Yacob that Medina-sarote
and Nunez were in love.
     There was from the first very great opposition to the marriage of Nunez
and Medina-sarote; not so much because they valued her as because they held
him as a being apart, an idiot, incompetent thing below the permissible
level of a man.  Her sisters opposed it bitterly as bringing discredit on
them all; and old Yacob, though he had formed a sort of liking for his
clumsy, obedient serf, shook his head and said the thing could not be.  The
young men were all angry at the idea of corrupting the race, and one went so
far as to revile and strike Nunez.  He struck back.  Then for the first time
he found an advantage in seeing, even by twilight, and after that fight was
over no one was disposed to raise a hand against him.  But they still found
his marriage impossible.
     Old Yacob had a tenderness for his last little daughter, and was
grieved to have her weep upon his shoulder.
     "You see, my dear, he's an idiot.  He has delusions; he can't do
anything right."
     "I know," wept Medina-sarote.  "But he's better than he was.  He's
getting better.  And he's strong, dear father, and kind–-stronger and kinder
than any other man in the world.  And he loves me–-and, father, I love him."
     Old Yacob was greatly distressed to find her inconsolable, and,
besides–-what made it more distressing–-he liked Nunez for many things.  So

                                                           Wells/BLIND/28
he went and sat in the windowless council-chamber with the other elders and
watched the trend of the talk, and said, at the proper time, "He's better
than he was.  Very likely, some day, we shall find him as sane as
ourselves."
     Then afterwards one of the elders, who thought deeply, had an idea.  He
was a great doctor among these people, their medicine-man, and he had a very
philosophical and inventive mind, and the idea of curing Nunez of his
peculiarities appealed to him.  One day when Yacob was present he returned
to the topic of Nunez.  "I have examined Nunez," he said, "and the case is
clearer to me.  I think very probably he might be cured."
     "This is what I have always hoped," said old Yacob.
     "His brain is affected," said the blind doctor.
     The elders murmured assent.
     "Now, WHAT affects it?"
     "Ah!"  said old Yacob.
     THIS," said the doctor, answering his own question.  "Those queer
things that are called the eyes, and which exist to make an agreeable
depression in the face, are diseased, in the case of Nunez, in such a way as
to affect his brain.  They are greatly distended, he has eyelashes, and his
eyelids move, and consequently his brain is in a state of constant
irritation and distraction."
     "Yes?"  said old Yacob.  "Yes?"
     "And I think I may say with reasonable certainty that, in order to cure
him complete, all that we need to do is a simple and easy surgical
operation–-namely, to remove these irritant bodies."

                                                           Wells/BLIND/29
     "And then he will be sane?"
     "Then he will be perfectly sane, and a quite admirable citizen."
     "Thank Heaven for science!"  said old Yacob, and went forth at once to
tell Nunez of his happy hopes.
     But Nunez's manner of receiving the good news struck him as being cold
and disappointing.
     "One might think," he said, "from the tone you take that you did not
care for my daughter."
     It was Medina-sarote who persuaded Nunez to face the blind surgeons.
     "YOU do not want me," he said, "to lose my gift of sight?"
     She shook her head.
     "My world is sight."
     Her head drooped lower.
     "There are the beautiful things, the beautiful little things–-the
flowers, the lichens amidst the rocks, the light and softness on a piece of
fur, the far sky with its drifting dawn of clouds, the sunsets and the
stars.  And there is YOU. For you alone it is good to have sight, to see
your sweet, serene face, your kindly lips, your dear, beautiful hands folded
together.  .  .  .  .  It is these eyes of mine you won, these eyes that
hold me to you, that these idiots seek.  Instead, I must touch you, hear
you, and never see you again.  I must come under that roof of rock and stone
and darkness, that horrible roof under which your imaginations stoop .  .  .
NO; YOU would not have me do that?"
     A disagreeable doubt had arisen in him.  He stopped and left the thing
a question.

                                                      Wells/BLIND/30
"I wish," she said, "sometimes–-" She paused.
"Yes?"  he said, a little apprehensively.
"I wish sometimes–-you would not talk like that."
"Like what?"
"I know it's pretty–-it's your imagination.  I love it, but NOW–-"
He felt cold.  "NOW?" he said, faintly.
She sat quite still.
"You mean–-you think–-I should be better, better perhaps–-"
He was realising things very swiftly.  He felt anger perhaps, anger at
the dull course of fate, but also sympathy for her lack of understanding–-a
sympathy near akin to pity.
"DEAR," he said, and he could see by her whiteness how tensely her
spirit pressed against the things she could not say.  He put his arms about
her, he kissed her ear, and they sat for a time in silence.
"If I were to consent to this?"  he said at last, in a voice that was
very gentle.
She flung her arms about him, weeping wildly.  "Oh, if you would," she
sobbed, "if only you would!"
For a week before the operation that was to raise him from his
servitude and inferiority to the level of a blind citizen Nunez knew nothing
of sleep, and all through the warm, sunlit hours, while the others slumbered
happily, he sat brooding or wandered aimlessly, trying to bring his mind to
bear on his dilemma.  He had given his answer, he had given his consent, and
still he was not sure.  And at last work-time was over, the sun rose in
splendour over the golden crests, and his last day of vision began for him.

                                                           Wells/BLIND/31
He had a few minutes with Medina-sarote before she went apart to sleep.
     "To-morrow," he said, "I shall see no more."
     "Dear heart!"  she answered, and pressed his hands with all her
strength.
     "They will hurt you but little," she said; "and you are going through
this pain, you are going through it, dear lover, for ME . .  .  .  Dear, if
a woman's heart and life can do it, I will repay you.  My dearest one, my
dearest with the tender voice, I will repay."
     He was drenched in pity for himself and her.
     He held her in his arms, and pressed his lips to hers and looked on her
sweet face for the last time.  "Good-bye!"  he whispered to that dear sight,
"good-bye!"
     And then in silence he turned away from her.
     She could hear his slow retreating footsteps, and something in the
rhythm of them threw her into a passion of weeping.
     He walked away.
     He had fully meant to go to a lonely place where the meadows were
beautiful with white narcissus, and there remain until the hour of his
sacrifice should come, but as he walked he lifted up his eyes and saw the
morning, the morning like an angel in golden armour, marching down the
steeps .  .  .  .
     It seemed to him that before this splendour he and this blind world in
the valley, and his love and all, were no more than a pit of sin.
     He did not turn aside as he had meant to do, but went on and passed
through the wall of the circumference and out upon the rocks, and his eyes

                                                           Wells/BLIND/32
were always upon the sunlit ice and snow.
     He saw their infinite beauty, and his imagination soared over them to
the things beyond he was now to resign for ever!
     He thought of that great free world that he was parted from, the world
that was his own, and he had a vision of those further slopes, distance
beyond distance, with Bogota, a place of multitudinous stirring beauty, a
glory by day, a luminous mystery by night, a place of palaces and fountains
and statues and white houses, lying beautifully in the middle distance.  He
thought how for a day or so one might come down through passes drawing ever
nearer and nearer to its busy streets and ways.  He thought of the river
journey, day by day, from great Bogota to the still vaster world beyond,
through towns and villages, forest and desert places, the rushing river day
by day, until its banks receded, and the big steamers came splashing by and
one had reached the sea–-the limitless sea, with its thousand islands, its
thousands of islands, and its ships seen dimly far away in their incessant
journeyings round and about that greater world.  And there, unpent by
mountains, one saw the sky–-the sky, not such a disc as one saw it here, but
an arch of immeasurable blue, a deep of deeps in which the circling stars
were floating .  .  .  .
     His eyes began to scrutinise the great curtain of the mountains with a
keener inquiry.
     For example; if one went so, up that gully and to that chimney there,
then one might come out high among those stunted pines that ran round in a
sort of shelf and rose still higher and higher as it passed above the gorge.
And then?  That talus might be managed.  Thence perhaps a climb might be

                                                           Wells/BLIND/33
found to take him up to the precipice that came below the snow; and if that
chimney failed, then another farther to the east might serve his purpose
better.  And then?  Then one would be out upon the amber-lit snow there, and
half-way up to the crest of those beautiful desolations.  And suppose one
had good fortune!
     He glanced back at the village, then turned right round and regarded it
with folded arms.
     He thought of Medina-sarote, and she had become small and remote.
     He turned again towards the mountain wall down which the day had come
to him.
     Then very circumspectly he began his climb.
     When sunset came he was not longer climbing, but he was far and high.
His clothes were torn, his limbs were bloodstained, he was bruised in many
places, but he lay as if he were at his ease, and there was a smile on his
face.
     From where he rested the valley seemed as if it were in a pit and
nearly a mile below.  Already it was dim with haze and shadow, though the
mountain summits around him were things of light and fire.  The mountain
summits around him were things of light and fire, and the little things in
the rocks near at hand were drenched with light and beauty, a vein of green
mineral piercing the grey, a flash of small crystal here and there, a
minute, minutely-beautiful orange lichen close beside his face.  There were
deep, mysterious shadows in the gorge, blue deepening into purple, and
purple into a luminous darkness, and overhead was the illimitable vastness
of the sky.  But he heeded these things no longer, but lay quite still

                                                           Wells/BLIND/34
there, smiling as if he were content now merely to have escaped from the
valley of the Blind, in which he had thought to be King.  And the glow of
the sunset passed, and the night came, and still he lay there, under the
cold, clear stars.
                                 # # # # #